# Enhancing Accuracy in Differentially Private Distributed Optimization Through Sensitivity Reduction

### v1

@@ 3.1 Algorithm Design
Consider two differentially private distributed algorithms (Alg 1 and Alg 2). Let $\Delta_{1}(k)$ and $\Delta_{2}(k)$ be the sensitivities of the two algorithms, respectively. Under Definition 1, the following claims hold:
@@
    Alg 1: agents use their noisy states in the consensus step. Alg 2: agents use their true states in the consensus step. Then, $\Delta_{1}(k)\leq\Delta_{2}(k)$ for all $k\geq 1$.
@@
    Alg 1: agents compute gradients based on noisy states. Alg 2: agents compute gradients based on true states. Then, $\Delta_{1}(k)\leq\Delta_{2}(k)$ for all $k\geq 1$.
@@
    Alg 1: agents exchange one variable $\bm{x}$. Alg 2: agents exchange two variables $\bm{x}$ and $\bm{y}$. Then, $\Delta_{1}(k)\leq\Delta_{2}(k)$ for all $k\geq 1$.
@@
Algorithm Closed-form expression Number of Communica-
for noise w.r.t. budget $\epsilon$ added noise tion
@@
Algorithm 1 involves four tunable parameters $\gamma$, $\beta$, $q_{1}$, and $q_{2}$. For these parameters, we provide two tuning strategies in Section 5. Using parameters (5), Algorithm 1 achieves $\epsilon$-DP, convergence, and a guaranteed level of accuracy. These properties will be analyzed in detail in Sections 4 and 5. Moreover, we summarize the advantages of Algorithm 1 over existing state-of-the-art algorithms [13, 15, 14, 16] in Table 1, with detailed discussions provided in the following remarks.
@@
For Algorithm 1, given any privacy budget $\epsilon>0$, the noise parameter $\nu_{k}$ can be easily obtained using Eq. (5). This is more convenient and practical compared to algorithms in [13, 15, 14, 16], in which there are no closed-form expression for the noise parameter with respect to the privacy budget.
@@
In the proposed algorithm, only variable $\bm{x}$ is shared among agents, while the auxiliary variable $\bm{y}$ remains local and is not exchanged. Compared to existing algorithms [13, 15, 14, 16], the proposed algorithm reduces both the number of communication variables and the required injected noise by half. As a result, it significantly reduces communication and computational burdens.
@@ 3.2 Related Methods
To get a better insight into the update rule of our algorithm, we compare it with the Distributed Gradient Descent (DGD) and Gradient Tracking (GT) methods. Notably, [12] and [14] develop differentially private distributed algorithms based on the (projected) DGD method. The update rule for DGD is
@@
The GT method overcomes the above limitation by introducing an auxiliary variable $\mathbf{y}$, which enables exact convergence with a constant stepsize. Note that the differentially private distributed algorithms in [13, 15, 14, 16] are based on the GT method or its variants. The update rule for GT is given by:
@@ 4 Differential Privacy
([12, Lemma 2]) If each agent $i$ adds a noise vector $\bm{\xi}_{i}(k)$ under Assumption 3 such that $\sum_{k=1}^{\infty}\frac{\Delta(k)}{\nu_{k}}\leq\epsilon$, then Algorithm 1 is $\epsilon$-differentially private.
In the following theorem, we first give an upper bound on the sensitivity of Algorithm 1. Then, based on Lemma 4.4, we prove that Algorithm 1, with the parameter settings given in (5), is $\epsilon$-differentially private.
@@
With the same initial state $\mathbf{x}(0)$ and observation sequence $\mathcal{Z}$, for any two adjacent problems $\mathcal{P}$ and $\mathcal{P}^{\prime}$, we have $\bm{z}_{i}(k)=\bm{z}_{i}^{\prime}(k)$ for all $i\in\mathcal{N}$ and all $k\geq 1$. Based on (4a), (4b), and $\bm{y}_{i}(0)=\bm{y}_{i}^{\prime}(0)$, it follows that $\bm{y}_{i}(k)=\bm{y}_{i}^{\prime}(k)$ for all $i\in\mathcal{N}$ and all $k\geq 1$. From Definition 1, we know that $f_{j}\not=f_{j}^{\prime}$ if and only if $j=i_{0}$. In light of (4c), it follows that $\bm{x}_{j}(k)=\bm{x}_{j}^{\prime}(k)$ for all $j\in\mathcal{N}$, $j\not=i_{0}$ and all $k\geq 1$. Therefore, the sensitivity of Algorithm 1 at iteration $k$ is bounded as follows:
@@
In Table 2, we compare the sensitivity of differentially private distributed algorithms under the same definitions and assumptions outlined in this paper. As shown, our algorithm demonstrates lower sensitivity compared to algorithms in [13, 15, 14, 16]. This advantage stems from its adherence to the design principles outlined in Proposition 1, whereas other algorithms did not consider the sensitivity. According to Lemma 4.4, for a fixed privacy budget, the algorithm’s sensitivity is positively correlated with the noise variance. That is, lower sensitivity requires smaller noise variance. This leads to possible improvement of the optimization accuracy. Therefore, under the same privacy budget, our algorithm can achieve higher optimization accuracy than existing state-of-the-art algorithms.
@@ 5.1 Preliminary Analysis and Fundamental Lemmas
([18, Lemma 2]) Under Assumption 1, the following inequalities hold:
@@ 5.2 Main Results
Before presenting the main results, we first state some useful equations and inequalities. From the parameter settings in (5), we obtain $\frac{\alpha_{k+1}}{\alpha_{k}}=q_{1}<1$ and $\nu_{k+1}\leq\nu_{k}$. According to Assumption 2 and Remark 5.8, we know that both $W$ and $\tilde{W}(k)$ are symmetric doubly-stochastic matrices. This implies that the eigenvalues of $W$ and $\tilde{W}(k)$ lie within $[-1,1]$. Since the induced graphs are connected, both $W$ and $\tilde{W}(k)$ have a unique largest eigenvalue equal to 1, with eigenvector $\bm{1}$. Thus, we have $\|W\|=1$, $\|\tilde{W}(k)\|=1$, and $\|\tilde{W}(k)-I\|\leq 2$. Since $10<\alpha_{k+1}\beta=\gamma\beta q_{1}^{k}<$ 1, we have $\|(1+\alpha_{k+1}\beta)W-\alpha_{k+1}\beta I\|\leq 3$ and $\|(1+\alpha_{k+1}\beta)W-\alpha_{k+1}\beta I-\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\|\leq 3$. In addition, $\|I-\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\|=1$.
Next, we present three inequalities in Lemmas 5.14, 5.15, and 5.16, which will assist in deriving our main results. The proofs of these lemmas are provided in Appendix 9.
@@
$\quad+2np\Big{(}15+6L\alpha_{k}+4L\alpha_{k+1}+npL\alpha_{k+1}+2L^{2}\alpha_{k}\alpha_{k+1}$
@@
$\quad+4np\Big{(}19+6L\alpha_{k}+5L\alpha_{k+1}+npL\alpha_{k+1}+2L^{2}\alpha_{k}\alpha_{k+1}$
@@
Let $s_{1}(k)\triangleq\mathbb{E}\left[\|\bar{\bm{x}}(k)-\bm{x}^{*}\|^{2}\right]$, $s_{2}(k)\triangleq\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}\right]$, and $s_{3}(k)\triangleq\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{x}(k-1)\|^{2}\right]$. In light of Lemmas 5.14, 5.15, and 5.16, we have
@@
For the case of $k<\infty$, we will demonstrate that $\rho(A_{k})$ is strictly less than 1. By Lemma 5.12, it suffices to ensure that $[A_{k}]_{11},[A_{k}]_{22},[A_{k}]_{33}<1$ and $\det(I-A_{k})>0$. If $\alpha_{k+1}\leq\frac{2}{\mu+L}$, then $10\leq[A_{k}]_{11}=1-\alpha_{k+1}\mu<$ 1. To guarantee that $[A_{k}]_{22}<1$ and $[A_{k}]_{33}<1$, we only need $\alpha_{k+1}^{2}<\frac{(1-\sigma^{2})^{2}}{12(3+\sigma^{2})L^{2}}$. If $\alpha_{k+1}$ further satisfies
@@
where $\tilde{\bm{s}}(k)\triangleq\left[s_{2}(k),s_{3}(k)\right]^{T}$, $\tilde{A}$ is defined as in (29), and $[\tilde{\bm{b}}_{k}]_{1}\triangleq\frac{2(2+\sigma^{2})(2nc_{1}+2c_{2}+c_{3})L^{2}\alpha_{k+1}^{2}}{1-\sigma^{2}}+[\bm{b}_{k}]_{2}$, $[\tilde{\bm{b}}_{k}]_{2}\triangleq\frac{(3+\sigma^{2})(4nc_{1}+4c_{2}+3c_{3})L^{2}\alpha_{k+1}^{2}}{1-\sigma^{2}}+[\bm{b}_{k}]_{3}$.
@@
where $c_{6}\triangleq\frac{\mu^{2}(\theta-1)(1-\sigma^{2})^{4}}{24(\theta+1)(\theta+2)(2+\sigma^{2})(3+\sigma^{2})L^{2}q_{1}^{4}}$. In summary, the stepsize $\alpha_{k}$ must satisfy conditions (26), (35), (38), (40), and $\alpha_{k+1}\leq\frac{2}{\mu+L}$. Since $\alpha_{k}$ decays exponentially, there always exists a $K$ such that $\alpha_{k}$ meets all the aforementioned conditions for all $k\geq K$. Let $k=k-K$. Considering (30) and (37), we complete the proof.
@@ 6 Numerical Results
In this section, we demonstrate the effectiveness of Algorithm 1 using the canonical sensor fusion problem. In this problem, all $n$ sensors collaborate to estimate an unknown variable $\bm{x}$ by solving the following optimization problem:
@@
where $\bm{v}_{i}\in\mathbb{R}^{m}$, $M_{i}\in\mathbb{R}^{m\times p}$, and $\omega_{i}\in\mathbb{R}$ represent the observation, measurement matrix, and regularization coefficient of sensor $i$, respectively. The parameter settings are consistent with those in [10].
Figure 1: The communication topology of 100 sensors.

### v2

@@ 3.1 Algorithm Design
Consider two differentially private distributed algorithms, denoted as Alg 1 and Alg 2, with sensitivities $\Delta_{1}(k)$ and $\Delta_{2}(k)$, respectively. Under the conditions given in Definition 1, the following claims hold:
@@
    Consensus step design: If Alg 1 has each agent use its own noisy state in the consensus step, while Alg 2 has each agent use its own true state, then $\Delta_{1}(k)\leq\Delta_{2}(k)$ for all $k\geq 1$.
@@
    Gradient computation design: If Alg 1 computes gradients based on noisy states while Alg 2 computes gradients based on true states, then $\Delta_{1}(k)\leq\Delta_{2}(k)$ for all $k\geq 1$.
@@
    Number of exchanged variables: If Alg 1 exchanges only variable $\bm{x}$ while Alg 2 exchanges two variables $\bm{x}$ and $\bm{y}$, then $\Delta_{1}(k)\leq\Delta_{2}(k)$ for all $k\geq 1$.
@@
Proposition 1 reveals that in differentially private distributed optimization, more extensive use of noisy information (i.e., using noisy states in both consensus and gradient computation) together with reducing the number of variables exchanged between agents can systematically reduce the algorithm’s sensitivity. The intuition is twofold. Firstly, using noisy information acts like a “blurring” process on the exact information, which mitigates the influence of an agent’s private data on its outputs, thereby reducing sensitivity. Secondly, reducing the shared information means an agent exposes less data externally, which naturally limits the amount of private information that could be disclosed and thus lowers sensitivity. Crucially, a lower sensitivity means that less noise is required to achieve the same privacy protection level. This directly translates into higher accuracy for the algorithm under an identical privacy budget.
@@
Algorithm Explicit, practical Number of Communica-
expression for $\nu_{k}(\epsilon)$ added noise tion
[20] $2p$ $2p$
@@
Algorithm 1 involves four tunable parameters $\gamma$, $\beta$, $q_{1}$, and $q_{2}$. For these parameters, we provide two tuning strategies in Section 5. Using parameters (5), Algorithm 1 achieves $\epsilon$-DP, convergence, and a guaranteed level of accuracy. These properties will be analyzed in detail in Sections 4 and 5. Moreover, we summarize the advantages of Algorithm 1 over existing state-of-the-art algorithms [20, 21, 23, 22, 24] in Table 2, with detailed discussions provided in the following remarks.
@@
For Algorithm 1, given any privacy budget $\epsilon>0$, the noise parameter $\nu_{k}$ can be directly and explicitly computed offline using the closed-form expression in Eq. (5). This provides a straightforward design rule, contrasting with the implicit or complex analytical relations in [20, 21, 23, 22, 24], which primarily serve for posterior analysis or as verification constraints rather than direct synthesis tools.
@@
In the proposed algorithm, only variable $\bm{x}$ is shared among agents, while the auxiliary variable $\bm{y}$ is kept local. Compared to existing GT-based algorithms [20, 23, 22, 24, 21], which require exchanging two variables, the proposed algorithm reduces both the number of communication variables and the amount of injected noise by half, significantly lowering communication and computational overhead. Note that the DPOP method [18] also exchanges only one variable. However, DPOP is based on DGD and relies on a decaying stepsize for convergence, which leads to a sublinear convergence rate even without privacy noise. In contrast, our algorithm employs an implicit gradient-tracking mechanism (see the next subsection for a detailed comparison among DGD, GT, and our scheme), which enables more effective gradient alignment. Under the same privacy budget, our algorithm achieves higher optimization accuracy (see Fig. 2 in Section 6), demonstrating a superior privacy-accuracy trade-off inherited from the GT framework.
@@ 3.2 Related Methods
To get a better insight into the update rule of our algorithm, we compare it with the DGD and GT methods. Notably, [18] and [22] develop differentially private distributed algorithms based on the DGD method. The update rule for DGD is
@@
The GT method overcomes the above limitation by introducing an auxiliary variable $\mathbf{y}$, which enables exact convergence with a constant stepsize. Note that the differentially private distributed algorithms in [20, 23, 22, 24, 21] are based on the GT method or its variants. The update rule for GT is given by:
@@ 4 Differential Privacy
([18, Lemma 2]) If each agent $i$ adds a noise vector $\bm{\xi}_{i}(k)$ under Assumption 3 such that $\sum_{k=1}^{\infty}\frac{\Delta(k)}{\nu_{k}}\leq\epsilon$, then Algorithm 1 is $\epsilon$-differentially private.
According to Lemma 4.4, ensuring that the algorithm achieves $\epsilon$-DP requires satisfying two conditions: (i) the algorithm’s sensitivity $\Delta(k)$ is bounded, and (ii) the variance $\nu_{k}$ of the injected noise is sufficiently large such that $\sum_{k=1}^{\infty}\frac{\Delta(k)}{\nu_{k}}\leq\epsilon$. Furthermore, it is established that the variance of the required noise is proportional to the sensitivity. This implies that the sensitivity bound directly determines the lower limit on the scale of noise that must be added to achieve $\epsilon$-DP. In the following theorem, we first give an upper bound on the sensitivity of Algorithm 1. Then, based on Lemma 4.4, we prove that Algorithm 1, with the parameter settings given in (5), is $\epsilon$-differentially private.
@@
With the same initial states $\mathbf{x}(0)$, $\mathbf{y}(0)$ and the same observation sequence $\mathcal{Z}$, for any two adjacent problems $\mathcal{P}$ and $\mathcal{P}^{\prime}$, we have $\bm{z}_{i}(k)=\bm{z}_{i}^{\prime}(k)$ for all $i\in\mathcal{N}$ and all $k\geq 1$. Based on (4a) and (4b), it follows that $\bm{y}_{i}(k)=\bm{y}_{i}^{\prime}(k)$ for all $i\in\mathcal{N}$ and all $k\geq 1$. From Definition 1, we know that $f_{j}\not=f_{j}^{\prime}$ if and only if $j=i_{0}$. In light of (4c), it follows that $\bm{x}_{j}(k)=\bm{x}_{j}^{\prime}(k)$ for all $j\in\mathcal{N}$, $j\not=i_{0}$ and all $k\geq 1$. Therefore, the sensitivity of Algorithm 1 at iteration $k$ is bounded by
@@
In Table 3, we compare algorithm sensitivities under the consistent framework. As Lemma 4.4 establishes, for a fixed privacy budget, sensitivity directly dictates the required noise variance: lower sensitivity enables smaller variance, which in turn improves optimization accuracy. As shown, our algorithm achieves the lowest sensitivity among the considered methods [20, 23, 22, 24, 21]. This stems from its adherence to the design principles outlined in Proposition 1, whereas other algorithms did not consider the sensitivity. Therefore, under the same privacy budget, our algorithm guarantees higher accuracy.
Remark 4.8.
In Theorem 4.5, the DP guarantee is stated for the infinite observation sequence $\mathcal{Z}$, but it directly applies to any adversary observing only a finite number of iterations $K$. This is because the $\epsilon$-DP guarantee on the entire sequence implies that the distributions of any finite prefix of the sequence are also $\epsilon$-indistinguishable. Therefore, our privacy guarantee is robust for practical (finite-observation) threat models.
@@ 5.1 Preliminary Analysis and Fundamental Lemmas
([33, Lemma 2]) Under Assumption 1, the following inequalities hold:
@@ 5.2 Main Results
Before presenting the main results, we first state some useful equations and inequalities. From the parameter settings in (5), we obtain $\frac{\alpha_{k+1}}{\alpha_{k}}=q_{1}<1$ and $\nu_{k+1}\leq\nu_{k}$. According to Assumption 2 and Remark 5.9, we know that both $W$ and $\tilde{W}(k)$ are symmetric doubly-stochastic matrices. This implies that the eigenvalues of $W$ and $\tilde{W}(k)$ lie within [1,1][-1,1]. Since the induced graphs are connected, both $W$ and $\tilde{W}(k)$ have a unique largest eigenvalue equal to 1, with eigenvector $\bm{1}$. Thus, we have $\|W\|=1$, $\|\tilde{W}(k)\|=1$, and $\|\tilde{W}(k)-I\|\leq 2$. Since $0<\alpha_{k+1}\beta=\gamma\beta q_{1}^{k}<1$, we have $\|(1+\alpha_{k+1}\beta)W-\alpha_{k+1}\beta I\|\leq 3$ and $\|(1+\alpha_{k+1}\beta)W-\alpha_{k+1}\beta I-\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\|\leq 3$. In addition, $\|I-\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\|=1$.
Next, we present three inequalities in Lemmas 5.15, 5.16, and 5.17, which will assist in deriving our main results. The proofs of these lemmas are provided in Appendix 9.
@@
[…]
@@
[…]
@@
Let $s_{1}(k)\triangleq\mathbb{E}\left[\|\bar{\bm{x}}(k)-\bm{x}^{*}\|^{2}\right]$, $s_{2}(k)\triangleq\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{1}\bar{\bm{x}}(k)\|^{2}\right]$, and $s_{3}(k)\triangleq\mathbb{E}\left[\|\mathbf{x}(k)-\mathbf{x}(k-1)\|^{2}\right]$. In light of Lemmas 5.15, 5.16, and 5.17, we have
@@
For the case of $k<\infty$, we demonstrate that $\rho(A_{k})$ is strictly less than 1. By Lemma 5.13, it suffices to ensure that $A_{k}]_{11},[A_{k}]_{22},[A_{k}]_{33}<1$ and $\det(I-A_{k})>0$. If $\alpha_{k+1}\leq\frac{2}{\mu+L}$, then $0\leq[A_{k}]_{11}=1-\alpha_{k+1}\mu<1$. To ensure $A_{k}]_{22}<1$ and $A_{k}]_{33}<1$, it is sufficient that $\alpha_{k+1}^{2}<\frac{(1-\sigma^{2})^{2}}{12(3+\sigma^{2})L^{2}}$. To satisfy the more stringent condition $\det(I-A_{k})>0$, a tighter bound on $\alpha_{k+1}$ is required. If $\alpha_{k+1}$ further satisfies
@@
where $\tilde{\bm{s}}(k)\triangleq\left[s_{2}(k),s_{3}(k)\right]^{T}$, $\tilde{A}$ is defined as in (29), and $\tilde{\bm{b}}_{k}]_{1}\triangleq\frac{2(2+\sigma^{2})(2nc_{1}+2c_{2}+c_{3})L^{2}\alpha_{k+1}^{2}}{1-\sigma^{2}}+[\bm{b}_{k}]_{2}$, $\tilde{\bm{b}}_{k}]_{2}\triangleq\frac{(3+\sigma^{2})(4nc_{1}+4c_{2}+3c_{3})L^{2}\alpha_{k+1}^{2}}{1-\sigma^{2}}+[\bm{b}_{k}]_{3}$.
@@
where $c_{6}\triangleq\frac{\mu^{2}(\theta-1)(1-\sigma^{2})^{4}}{24(\theta+1)(\theta+2)(2+\sigma^{2})(3+\sigma^{2})L^{2}q_{1}^{4}}$. In summary, the stepsize $\alpha_{k}$ must satisfy conditions (26), (35), (38), (40), and $\alpha_{k+1}\leq\frac{2}{\mu+L}$. Since $\alpha_{k}$ decays exponentially, there always exists a $K$ such that $\alpha_{k}$ meets all the aforementioned conditions for all $k\geq K$. Let k=kKk=k-K. Considering (30) and (37), we complete the proof.
@@ 6 Numerical Results
Figure 1: The communication topology of 100 sensors.
Figure 2: Comparison of convergence between Algorithm 1 and the DPOP method [18] under the same privacy budget in distributed estimation problems.
Figure 3: Comparison of convergence between Algorithm 1 and the PP-DOAGT method [24] under the same noise in distributed estimation problems.
Figure 4: Comparison of convergence between Algorithm 1 and the Diff-TV-AB method [21] under the same noise in distributed estimation problems.
6.1 Comparison of Differentially Private Methods Based on Distributed Sensor Fusion
In this subsection, we validate the effectiveness of Algorithm 1 by comparing it with existing differentially private methods on the canonical distributed sensor fusion problem. In this problem, all $n$ sensors collaborate to estimate an unknown variable $\bm{x}$ by solving the following optimization problem:
@@ 6.1 Comparison of Differentially Private Methods Based on Distributed Sensor Fusion
where $\bm{v}_{i}\in\mathbb{R}^{m}$, $M_{i}\in\mathbb{R}^{m\times p}$, and $\omega_{i}\in\mathbb{R}$ represent the observation, measurement matrix, and regularization coefficient of sensor $i$, respectively. The parameter settings are consistent with those in [15]. We consider a multi-agent system of 100 sensors, where the communication network is generated using an Erdős-Rényi (ER) random graph model with an edge creation probability of 0.1, as illustrated in Fig. 1. The dimensions in Problem (42) are set as $m=3$ and $p=2$.
For Algorithm 1, the initial states $\bm{x}_{i}(0)$ are randomly sampled from a standard normal distribution. The stepsize $\alpha_{k}$ and noise parameter $\nu_{k}$ are set according to (5), and the relevant parameters $\gamma$, $\beta$, $q_{2}$, and $q_{1}$ are tuned following the heuristic strategy described in Remark 5.20. We implement Algorithm 1 under different privacy budgets ($\epsilon\in\{0.1,1,10\}$). For $\epsilon=0.1/1/10$, the parameters are set as $\gamma=0.001/0.001/0.002$, $\beta=1000/1000/100$, $q_{2}=0.99/0.99/0.99$, and $q_{1}=0.92/0.97/0.97$, respectively. For each case, we calculate the mean and standard deviation of the residuals (denoted as $\|\mathbf{x}(k)-\mathbf{x}^{*}\|^{2}$}) over 1000 trials, with the results shown as the red line in Fig. 2. It can be observed that our algorithm quickly approaches the optimal solution under all privacy budgets. This confirms that the proposed algorithm can achieve arbitrary $\epsilon$-DP. Moreover, higher privacy levels result in lower optimization accuracy, which aligns with the conclusion of Theorem 5.18.
For comparison, we implement the DPOP method from [18] under the same privacy budgets. The corresponding results are shown by the blue lines in Fig. 2. It can be seen that our algorithm achieves higher accuracy than the DPOP method. This is because Algorithm 1’s sensitivity is more compact, and the introduction of auxiliary variable $\bm{y}_{i}$ enhances its convergence. Furthermore, we compare the proposed algorithm with the latest PP-DOAGT method [24] and the Diff-TV-AB method [21]. Note that, unlike our algorithm and DPOP, both PP-DOAGT and Diff-TV-AB do not provide an explicit closed-form expression that directly maps a given privacy budget to the required noise parameters. Instead, they offer an implicit and complex relationship in which the privacy budget is expressed as a function of the noise and other internal parameters (see Eq. (17) or Eq. (22) in [24] and Theorem 3 in [21]). Such a relationship cannot be directly used to derive deployable noise sequences under a prescribed privacy budget. Thus, to enable a meaningful comparison, we inject the same noise-variance sequence (i.e., the same noise conditions) into all three algorithms. In addition, since our work does not require Assumption 3 of [21] to achieve DP, for fairness, we replace the constant stepsize in Diff-TV-AB with a diminishing stepsize. The results are shown in Figs. 3 and 4. Under higher privacy levels, with privacy budgets computed via the theoretical formulas of PP-DOAGT itself being 18446.44, 3652.67, and 365.27 (where $\delta=1$), and those computed via the theoretical formulas of Diff-TV-AB itself being 102.40, 20.34, and 7.12 (where $\vartheta=1$), our algorithm achieves optimization accuracy comparable to that of the PP-DOAGT and Diff-TV-AB methods. This is because our algorithm exhibits lower sensitivity than PP-DOAGT and Diff-TV-AB, as shown in Table 3.
Figure 5: Running time and communication cost comparison of different algorithms.
To further validate the practical efficiency of Algorithm 1, we conduct a comparative experiment among four methods: Algorithm 1, DPOP, PP-DOAGT, and Diff-TV-AB under identical problem configurations. We adopt the same stopping criterion, that is, the iteration is terminated when the optimization residual falls below $1\times 10^{-2}$. To ensure statistical reliability, we perform 100 valid independent runs for each method and record the running time and total communication cost for every trial. The results are presented in Fig. 5. Algorithm 1 achieves the lowest median running time and the smallest median communication cost, demonstrating significantly better performance than PP-DOAGT and Diff-TV-AB. This advantage can be attributed to two main factors. On one hand, compared with DPOP, Algorithm 1 exhibits better convergence behavior, reaching the prescribed accuracy in fewer iterations. On the other hand, relative to PP-DOAGT and Diff-TV-AB, Algorithm 1 significantly reduces the per-iteration computational and communication overhead (see Table 2), which yields higher efficiency in each iteration step.
6.2 Comparison of Differentially Private Methods Based on Distributed Learning on MNIST
In this subsection, we further validate the effectiveness of Algorithm 1 through distributed learning tasks on the MNIST dataset, while comprehensively evaluating its performance via parameter sensitivity analysis and privacy-accuracy trade-off studies. The MNIST dataset, a classic benchmark comprising handwritten digit images, is widely utilized for evaluating algorithm performance in image classification tasks. In this experiment, we consider a distributed learning scenario with 100 agents collaboratively training a convolutional neural network (CNN), whose communication topology is illustrated in Fig. 1. The CNN architecture adopts a two-layer convolutional structure: each layer consists of a $5\times 5$ convolution, BatchNorm normalization, and a $2\times 2$ max-pooling operation, followed by a fully-connected layer that outputs predictions over 10 classes. Each agent maintains a local copy of this CNN. We uniformly distribute the MNIST training images across all agents and train the model using a batch size of 16 with a cross-entropy loss function.